OpenStudy (31356):

Double-check this l^2-9l=l(l-9) {those are the letter l) Or is it 1(1-9)? MEDAL REWARDED!!

OpenStudy (31356):

@jim_thompson5910 @Mertsj

OpenStudy (31356):

@mathmale

OpenStudy (unklerhaukus):

\[\ell^2-9\ell=\ell\times\ell-9\times\ell=\ell(\ell-9)\] yes

OpenStudy (31356):

Thanks!